OpenStudy (anonymous):

which expression defines the given series for seven terms? 8+13+18+...

OpenStudy (anonymous):

A: (2n+6) B: (3n+5) C: (n+8) D: (5n+3)

OpenStudy (raden):

arithmetic series : Sn = n/2 (2*a1 + (n-1)d)

OpenStudy (anonymous):

what the letter d mean

OpenStudy (raden):

d = difference but looks your options above asking for just the sequence, not for series

OpenStudy (raden):

therefore use the formula : an = a1 + (n-1)d

OpenStudy (raden):

known a1 = 8, and d = a2 - a1 = 13 - 8 = 5 so, an = a1 + (n-1)d an = 8 + (n-1)5 an = 8 + 5n - 5 an = 5n + 3

OpenStudy (anonymous):

thanks

OpenStudy (raden):

welcome